Half of isha's age added to twice of majue's age total 70 years.in 5 years time their age added up to 75 years how old are they.
Answer by fractalier(6550)

Call isha's age and majue's age NOW, x and y. Then we can write
(1/2)x + 2y = 70 (let's multiply by two) and get
x + 4y = 140
In 5 years,
(x+5) + (y+5) = 75 or
x + y + 10 = 75 and
x + y = 65
Subtract this from the above and get
x + 4y = 140
-(x + y = 65)
---------------
3y = 75
y = 25
and then
x = 40